### v1

# EMLoC: Emulator-based Memory-efficient Fine-tuning with LoRA Correction

Abstract
Open-source foundation models have seen rapid adoption and development, enabling powerful general-purpose capabilities across diverse domains. However, fine-tuning large foundation models for domain-specific or personalized tasks remains prohibitively expensive for most users due to the significant memory overhead beyond that of inference. We introduce EMLoC, an Emulator-based Memory-efficient fine-tuning framework with LoRA Correction, which enables model fine-tuning within the same memory budget required for inference. EMLoC constructs a task-specific light-weight emulator using activation-aware singular value decomposition (SVD) on a small downstream calibration set. Fine-tuning then is performed on this lightweight emulator via LoRA. To tackle the misalignment between the original model and the compressed emulator, we propose a novel compensation algorithm to correct the fine-tuned LoRA module, which thus can be merged into the original model for inference. EMLoC supports flexible compression ratios and standard training pipelines, making it adaptable to a wide range of applications. Extensive experiments demonstrate that EMLoC outperforms other baselines across multiple datasets and modalities. Moreover, without quantization, EMLoC enables fine-tuning of a 38B model on a single 24GB consumer GPU—bringing efficient and practical model adaptation to individual users.
1 Introduction
General-purpose foundation models have demonstrated impressive zero-shot capabilities across a wide range of benchmarks [11, 8, 2, 35, 6]. For real-world deployment, such as domain-specific tasks [21, 42] or personalized user behavior [26, 30] , further customization by fine-tuning is still required. However, fine-tuning typically incurs significantly more memory overhead than inference [28]. Consequently, if users have a fixed amount of available computing resources, they will be forced to choose between two unfavorable options. First, they can use a small model that fits within their memory budget for fine-tuning as shown in Fig. 1(a), but this sacrifices the emergent capabilities [38] of larger models and underutilizes hardware during inference. Alternatively, they can opt for a large model that fully utilizes resources during inference but exceeds memory limits for fine-tuning as shown in Fig. 1(b), making user-specific adaptation infeasible and potentially limiting performance in specialized applications. This paper addresses a central research question: Is it possible to design a fine-tuning strategy such that users can fine-tune a model under the same memory budget as inference?
The memory cost of fine-tuning can be broadly attributed to three components: optimizer states, intermediate activations, and the model parameters themselves, as marked in Fig. 1 with different colors. Initial efforts to reduce the memory usage of fine-tuning concentrated on the first two components. The first component, optimizer states, stores auxiliary information such as momentum and variance in the Adam optimizer [15] for each trainable parameter. This overhead can be mitigated using parameter-efficient fine-tuning (PEFT) methods [13, 20, 41], such as LoRA [13], which reduces trainable parameters by only updating lightweight low-rank matrices. The second component, intermediate activations, refers to the values retained during the forward pass for use in backpropagation. Previous works reduced it by modifying model architectures [34, 10, 19] or gradient checkpointing [4]. However, since the components reduced by both strategies still presented during fine-tuning, reducing them alone, without addressing the memory usage of model parameters, cannot fully close the memory gap between fine-tuning and inference (Fig. 1(b)).
Few recent works have taken initial steps toward addressing the third component, the memory cost of model parameters. For example, Offsite-Tuning [40] proposes to drop some of the intermediate layers entirely during fine-tuning and to update only the top and bottom layers, thereby saving some memory occupied by model parameters. LORAM [45] introduces row-pruning as a method to reduce model parameters during fine-tuning, allowing the updating of the unpruned rows in the intermediate layers. Despite their promise, these methods typically rely on memory-intensive full-model training. This entails a strong and often impractical assumption that well-resourced institutions must be involved, making them unfeasible for individual users. In addition, they impose strict constraints on which subsets of weights can be updated, limiting their flexibility and general applicability across tasks and architectures. Moreover, they overlook the fact that different base models are used during fine-tuning and inference, introducing potential mismatches that degrade performance.
Figure 1: The dilemma caused by additional memory overhead during fine-tuning. (a) Users opt for a smaller 8B model, sacrificing emergent capabilities and underutilizing available hardware. (b) Use of a larger 26B model requiring memory exceeding the hardware limit even with LoRA [13] and gradient checkpointing [4] techniques. (c) Our EMLoC utilizes a smaller model during fine-tuning, allowing the same budget for both training and inference.
In this paper, we propose EMLoC, an Emulator-based Memory-efficient fine-tuning framework with LoRA Correction. EMLoC completely eliminates the memory cost gap between inference and training and can be carried out solely by individual users. The central idea is to perform fine-tuning on a low-rank model which we call the emulator so that the memory taken up by model parameters can be reduced as depicted in Fig. 1(c). To make this approach effective, we introduce an efficient, downstream-aware emulator construction procedure based on activation-aware singular value decomposition (SVD), using a small subset of downstream data for calibration. This yields a lightweight emulator tailored to the downstream data, without requiring costly full-model training. Fine-tuning is then conducted entirely on the emulator without any assumptions about the fine-tuning process—any standard training pipeline can be used. However, since the LoRA modules are fine-tuned on the emulator rather than the original model, naively merging the LoRA modules into the original model will suffer from the misalignment between the full model and the compressed emulator. To mitigate this issue, we further introduce a novel LoRA correction algorithm, which adjusts the parameters to compensate such discrepancy during inference.
Our contributions can be summarized as follows:
  • We introduce a pipeline that closes the memory cost gap between inference and fine-tuning, enabling users to fine-tune a model under the same memory budge as inference.
  • We propose a novel emulator-based fine-tuning framework that significantly reduces memory consumption by operating on a downstream-aware low-rank approximation of the model.
  • We develop a LoRA correction algorithm that addresses the misalignment between emulator-based fine-tuning and full-model inference, improving final performance and transferability.
2 Related Works
2.1 Reducing Optimizer State
There are mainly two ways to reduce the memory required for optimizer states: by reducing the number of trainable parameters [13, 41, 43, 46, 14, 12, 20], or by reducing auxiliary information required per parameter [31, 18, 49]. A prominent example of the first approach is LoRA [13], which introduces trainable low-rank adapters into frozen pre-trained models to achieve parameter-efficient fine-tuning. For the second approach, Adafactor [31] eliminates the need to store full-rank second-moment estimators by using a factored approximation. While these methods are effective, a non-trivial number of trainable parameters are required to maintain acceptable performance. As a result, the optimizer states associated with those parameters continue to contribute a non-negligible portion of the overall memory footprint.
2.2 Reducing Intermediate Activation
Another way to fine-tune with less memory overhead is to reduce memory from intermediate activations. Two explored strategies are shortening the backpropagation path [34, 25, 41, 44] and recomputing activations during the backward pass [4, 10, 19]. An example of the first strategy is Ladder Side-Tuning [34], where trainable modules are placed outside the backbone to avoid backpropagation through the backbone. [19] carefully design the position and initialization of the adapter modules so that the pre-trained model becomes reversible, which enables recomputing intermediate activation from the output during backpropagation. However, these methods cannot completely eliminate activation memory due to the fundamental need for gradients, and they typically incur additional computational overhead.
2.3 Reducing Model Parameters
Beyond optimizing optimizer states and intermediate activations, very recently reducing the memory footprint of the model parameters during fine-tuning has been explored. It is worth noting that quantization-based methods [9, 48] are largely orthogonal to our focus: they do not reduce fine-tuning memory relative to inference, and can therefore be combined with our method. An approach to reducing parameter memory during fine-tuning involves using a pruned model for training while maintaining a full model for inference [40, 45]. For example, LORAM [45] assumes that the model publisher provides a row-pruned and continually pre-trained variant of the original model. Users fine-tune this reduced model and then transfer the learned LoRA modules back to the full model. However, their setting requires continual pre-training of the pruned model, making it inaccessible to individual users. Furthermore, only the unpruned weights can be fine-tuned, limiting flexibility in selecting which parameters to adapt. In contrast, our work introduces a pipeline that can be performed by users individually and allows flexible selection of tunable parameters.
3 Method
Our proposed method, EMLoC, is a memory-efficient fine-tuning framework that enables individual users to fine-tune a model with the same memory cost as inference. An overview of EMLoC is shown in Fig. 2. EMLoC consists of three main stages. First, we construct a downstream-aware lightweight emulator by performing activation-aware singular value decomposition (SVD) [37] on a small calibration set drawn from the downstream data. Second, we fine-tune the emulator with LoRA modules using any standard training pipeline. The reduced parameter size of the emulator is crucial for enabling memory-efficient fine-tuning. Finally, to address the misalignment issue between the original model and the emulator, we present a novel LoRA correction algorithm to correct the learned LoRA modules, and transfer them back to the original model before inference.
Figure 2: Overview of EMLoC. Stage 1: Construct a downstream-aware lightweight emulator. Stage 2: Fine-tune the emulator via LoRA, allowing reduced memory costs. Stage 3: Update the LoRA module to compensate the misalignment between the full model and emulator.
3.1 Memory-Efficient Fine-Tuning: Problem Formulation
Given a pre-trained model $\mathcal{M}$ and a downstream dataset $\mathcal{D}_{\text{down}}$, the goal is to fine-tune $\mathcal{M}$ with the same memory constraint as inference. We denote the collection of linear weights in $\mathcal{M}$ as $\mathbb{W}=\{W_{i}\}$. In EMLoC, we introduce an lightweight emulator, denoted as $\mathcal{E}$, constructed by replacing each $W_{i}\in\mathbb{W}$ with its counterpart $W^{\mathcal{E}}_{i}\in\mathbb{W}^{\mathcal{E}}$. The subscript $i$ here is retained to emphasize the positional correspondence between $W_{i}$ and $W^{\mathcal{E}}_{i}$ in the architecture. For simpleness, we may omit the index $i$ when the context is clear. Fine-tuning is then performed on the emulator $\mathcal{E}$ using any standard training pipeline. During training, a set of LoRA modules $\mathbf{\Lambda}=\{\Lambda_{i}\}$ is learned on top of $W^{\mathcal{E}}$, resulting in the fine-tuned model weights $W^{\mathcal{E}}+\Lambda$. At inference time, the learned LoRA modules are transferred back to the original model $\mathcal{M}$, with the corresponding weights $W+\Lambda$, which is used for inference. Note that conventional parameter-efficient fine-tuning (PEFT) methods are a special case of this framework, where $\mathbb{W}^{\mathcal{E}}=\mathbb{W}$ and hence $\mathcal{E}=\mathcal{M}$.
3.2 Downstream-aware Emulator Construction and Fine-tuning
Downstream-aware Emulator Construction.
The goal of emulator construction is to create a lightweight replacement $\mathcal{E}$ for the original model $\mathcal{M}$ to be used during fine-tuning. To ensure that EMLoC is memory-efficient, generalizable, and effective, the emulator must satisfy three key criteria. First, it should have fewer parameters than the original model to reduce memory consumption during fine-tuning. Second, it should support flexible placement of LoRA modules, enabling all weights to be fine-tunable. In other words, if a LoRA module $\Lambda$ can be used to fine-tune a weight matrix $W$ in the original model, then EMLoC must support placing and training $\Lambda$ beside the corresponding $W^{\mathcal{E}}$ in the emulator, thereby enabling effective fine-tuning of $W$. Third, it should preserve knowledge relevant to the downstream task to ensure the fine-tuning process is effective. To meet all three criteria simultaneously, we propose constructing the emulator using activation-aware SVD, which naturally balances parameter reduction, flexibility, and task relevance.
Specifically, we construct the set of emulator weights $\mathbb{W}^{\mathcal{E}}$ by applying the activation-aware SVD method proposed in SVD-LLM [37] to each weight matrix $W\in\mathbb{W}$
$\mathbb{W}^{\mathcal{E}}=\left\{W^{\mathcal{E}}=W_{U}W_{V}=\text{SVD-LLM}(W,n)\mid W\in\mathbb{W}\right\}\text{,}$ (1)
where $n$ is the number of kept singular values, chosen to control the parameter count of the emulator. The emulator $\mathcal{E}$ is then constructed by replacing each $W$ in $\mathcal{M}$ with its corresponding low-rank approximation $W^{\mathcal{E}}$. Although we denote each $W^{\mathcal{E}}$ as a single matrix, it is stored in its factored form $W_{U}W_{V}$, which significantly reduces the memory footprint of model parameters, thus satisfying the first criterion. Importantly, each $W^{\mathcal{E}}$ retains the same position, and the input/output dimensions as $W$, differing only in the rank of the transformation. As a result, any LoRA module $\Lambda$ that could be applied to $W$ can equivalently be applied to $W^{\mathcal{E}}$, enabling all weights to be fine-tunable, the second criterion. Finally, SVD-LLM minimizes the output reconstruction error
$\|X^{\top}W-X^{\top}W^{\mathcal{E}}\|_{F},$ (2)
where $X$ is the intermediate activation computed from calibration data $\mathcal{D}_{C}$. This ensures that the emulator preserves task-relevant knowledge, making fine-tuning on it effective, thereby fulfilling the third criterion.
Fine-tuning.
Owing to the specific design choices in the emulator, we can apply any standard PEFT techniques [13, 41, 43] to fine-tune the significantly smaller emulator, achieving memory-efficient fine-tuning. In particular, if we aim to fine-tune a weight matrix $W_{i}$ in the original model, we instead insert a LoRA module next to $W^{\mathcal{E}}_{i}$ to fine-tune the emulator as in conventional frameworks. This allows us to use any existing training pipeline without modification, making our method a plug-and-play replacement for conventional LoRA fine-tuning, while further reducing memory usage.
Figure 3: LoRA correction to compensate model misalignment. To alleviate the misalignment that arises from fine-tuning the lightweight emulator but running inference on the original model, LoRA parameters are corrected via feature spaces between the emulator and the original model.
3.3 LoRA Correction for Compression Misalignment
Algorithm 1 LoRA_correction
Input: $W_{A},W_{B}$: LoRA weights;
$W,W^{\mathcal{E}}$: Linear weights from $\mathcal{M}$ and $\mathcal{E}$;
$\lambda$: An hyperparameter for regularization.
Output: $W^{c}_{A},W^{c}_{B}$: Corrected LoRA.
1. Apply SVD to $W_{A}$, obtaining $U,\Sigma,V^{T}$
2. $W^{\prime}_{A}\leftarrow U$, $W^{\prime}_{B}\leftarrow\Sigma V^{T}W_{B}$
3. $\Delta\leftarrow W^{\prime\top}_{A}(W-W^{\mathcal{E}})$
4. $W^{c}_{B}\leftarrow W^{\prime}_{A}$
5. $W^{c}_{B}\leftarrow W^{\prime}_{B}-\text{clamp}(\Delta,\lambda)$
return $W^{c}_{A},W^{c}_{B}$
To fully utilize the capacity of the original model during inference, we aim to transfer the LoRA modules $\Lambda$ fine-tuned on $\mathcal{E}$ back to the original model $\mathcal{M}$. However, the emulator construction step that replaces $W$ with its low-rank counterpart $W^{\mathcal{E}}$ inevitably introduces a misalignment between these two models. As a result, the base model outputs seen by the LoRA modules are different during fine-tuning and inference. The difference is unexpected by the LoRA modules since they were optimized in the context of $W^{\mathcal{E}}$. To address this issue, we propose a novel LoRA correction algorithm and correct the misalignment explicitly, as shown in Fig. 3.
Objective of LoRA correction.
The goal of our LoRA correction procedure is to correct the LoRA module $\Lambda$ fine-tuned with the emulator $\mathcal{E}$, and derive a new LoRA module $\Lambda^{c}$ suitable for the original model $\mathcal{M}$. In other word, for each input $x$ that activates the LoRA, the merged output should remain consistent between training and inference. Formally, we enforce the following condition:
$x^{\top}(W+\Lambda^{c})=x^{\top}(W^{\mathcal{E}}+\Lambda)\quad\forall x\in\mathcal{V}_{\Lambda}\text{,}$ (3)
where $\mathcal{V}_{\Lambda}$ denotes the subspace on which $\Lambda$ produces none-zero outputs—i.e., the region where it is active. It is worth noting that we enforce Eq. 3 only within $\mathcal{V}_{\Lambda}$, so that the correction has effects exclusively when the LoRA module is active. This design prevents unintended side effects in parts of the model unrelated to the LoRA module when doing correction, thereby preserving the integrity of the original model’s behavior.
Rewriting Eq. 3, we obtain an equivalent condition,
$x^{\top}(\Lambda^{c}-\Lambda)=x^{\top}(W-W^{\mathcal{E}})=\Delta_{x}\quad\forall x\in\mathcal{V}_{\Lambda}\text{.}$ (4)
This formulation highlights the explicit correction objective. The desired $\Lambda^{c}$ aims to compensate for the misalignment $\Delta_{x}$, which exists between the outputs of $W$ and $W^{\mathcal{E}}$. This can be achieved through offsetting the original LoRA module by the correction term $\Delta_{x}$.
Mitigating compression misalignment.
Although this idea is conceptually straightforward, two practical challenges arise. First, the condition in Eq. 4 must hold for all inputs $x\in\mathcal{V}_{\Lambda}$. Second, there is no direct way to incorporate the correction term $\Delta_{x}$ into the LoRA weights $\Lambda=W_{A}W_{B}$.
The first challenge is resolved by leveraging two facts: the condition in Eq. 4 is linear, and $\mathcal{V}_{\Lambda}$ forms a subspace of the input feature space. As a result, it suffices to enforce the condition only on a basis of $\mathcal{V}_{\Lambda}$, denoted by $\beta_{\Lambda}$
$x^{\top}(\Lambda^{c}-\Lambda)=x^{\top}(W-W^{\mathcal{E}})=\Delta_{x}\quad\forall x\in\beta_{\Lambda}\text{.}$ (5)
Furthermore, a natural choice for $\beta_{\Lambda}$ is the set of column vectors of $W_{A}$, the input projection matrix of the original LoRA module. To address the second challenge, we introduce a preprocessing step that reparameterizes the LoRA module while preserving its functionality. Specifically, we perform SVD on the input projection $W_{A}$, yielding $W_{A}=U\Sigma V^{\top}$. We then redefine the LoRA factors as
$W^{\prime}_{A}=U,\ \ \ W^{\prime}_{B}=\Sigma V^{\top}W_{B}\text{,}$ (6)
such that the overall transformation remains unchanged $\Lambda=W_{A}W_{B}=W^{\prime}_{A}W^{\prime}_{B}$. After this reparameterization, the columns of $W^{\prime}_{A}$, denoted by $\{a_{1},\ldots,a_{r}\}$, are orthonormal. This orthogonality makes the behavior of the LoRA module on each basis vector $a_{i}$ governed directly by the corresponding $b_{i}$, $i$-th row of $W^{\prime}_{B}$.
$a_{i}^{\top}\Lambda=\left(a_{i}^{\top}W^{\prime}_{A}\right)W^{\prime}_{B}=e_{i}^{\top}W^{\prime}_{B}=b_{i},$ (7)
where $e_{i}$ is the standard basis vector. In other words, modifying the output of $\Lambda$ on each basis vector $a_{i}$ reduces to adjusting the corresponding row $b_{i}$ in $W^{\prime}_{B}$.
With the challenges addressed, we now formally describe the LoRA correction algorithm, as presented in Algorithm 1. We begin by preprocessing the original LoRA module $\Lambda=W_{A}W_{B}$ to obtain $W^{\prime}_{A}$ and $W^{\prime}_{B}$. Next, we select the columns of $W^{\prime}_{A}$ as the basis $\beta_{\Lambda}$ in Eq. 3. The correction term $\Delta$ is then computed by measuring the output difference between the original weight $W$ and the emulator weight $W^{\mathcal{E}}$ under this basis
$\Delta=W^{\prime\top}_{A}\left(W-W^{\mathcal{E}}\right)\text{.}$ (8)
Each row of $\Delta$ represents the correction $\Delta_{a_{i}}$ corresponding to a basis vector $a_{i}$. Finally, we adjust $W^{\prime}_{B}$ to satisfy the condition in Eq. 5, compensating for the misalignment between $W$ and $W^{\mathcal{E}}$
$W^{c}_{A}=W^{\prime}_{A},\ \ \ W^{c}_{B}=W^{\prime}_{B}-\text{clamp}(\Delta,\lambda)\text{,}$ (9)
where the clamp function limits the norm of each correction vector $\Delta_{a_{i}}$ to be no more than a multiple $\lambda$ of the corresponding $b_{i}$. This hyperparameter $\lambda$ acts as a safeguard against excessive deviation, and we find it improves robustness in practice. The resulting corrected LoRA module, $\Lambda^{c}=W^{c}_{A}W^{c}_{B}$, is thus adapted to account for the misalignment introduced by training on the emulator and inferring on the original model.
Table 1: Performance comparisons of finetuning approaches on VQA with different memory costs. With InternVL2.5-8B as the original model, and 50% indicates finetuning on 4B models is conducted. Note that direct finetuning on InternVL2.5-8B (i.e., rows in gray) exceeds the memory budgets of all other finetuning methods and thus is not feasible under the same constraints.
Ratio Method ChartQA DocVQA InfoVQA TextVQA PMC-VQA WebSRC WC-VQA
Original 84.5 92.2 69.8 79.6 52.9 87.4 53.4
InternVL 4B 82.9 90.4 64.6 75.8 50.6 84.8 48.6
Offsite [40] 84.3 91.3 69.0 77.7 51.0 76.1 45.9
UPop [32] 84.4 92.0 69.7 78.5 50.7 76.4 42.1
50% EMLoC 84.6 92.3 69.9 78.8 52.3 85.2 48.8
Original 84.5 92.2 69.8 79.6 52.9 87.4 53.4
InternVL 2B 78.8 87.4 56.0 73.2 44.6 78.1 34.4
Offsite [40] 84.0 92.3 69.6 77.2 50.6 76.6 45.4
UPop [32] 84.3 92.2 69.6 78.2 50.9 76.6 44.1
25% EMLoC 84.2 92.3 70.0 79.0 51.6 79.6 46.2
Original 84.8 92.0 68.0 79.5 52.3 84.8 51.9
InternVL 2B 78.2 85.7 53.5 71.7 44.2 77.4 33.0
Offsite [40] 83.9 91.8 67.0 77.1 50.1 75.2 44.2
UPop [32] 84.2 91.9 67.1 78.1 50.3 75.8 43.9
25%+ QLoRA EMLoC 84.2 91.9 67.5 79.0 50.9 78.1 44.8
4 Experiment
4.1 Experiment Setup
Datasets.
In our main experiments, we evaluate EMLoC on seven visual question answering (VQA) datasets. Four of these are widely used benchmarks: ChartQA [22], DocVQA [23], InfoVQA [24], and TextVQA [33]. To reflect realistic deployment scenarios that require domain adaptation, we also include three datasets from specialized domains: PMC-VQA [47] (medical imaging), WebSRC [5] (web page understanding), and WC-VQA [39] (multicultural knowledge). Ablation studies and additional analyses are primarily conducted on these three domain-specific datasets. For language modeling, we evaluate on two challenging mathematical reasoning benchmarks: MathQA [1] and GSM8K [7]. All reported metrics across datasets are based on accuracy.
Comparison.
For the main experiments, we compare EMLoC against four baselines. Original: We directly fine-tune the original large model, InternVL2.5-8B [6], with LoRA. This baseline incurs significantly higher memory cost compared to other methods and serves as an upper bound in performance. Small: We fine-tune smaller variants of InternVL (4B and 2B), representing the low-resource alternative illustrated in Fig. 1(a). Offsite: We follow the Offsite-tuning [40] framework using the variant that omits full-model distillation to ensure fair comparison under the same compute constraints. UPop: A row-pruning-based method [32] is used to construct the emulator, followed by LoRA transfer using the procedure described in LORAM [45]. Except for the Small baseline, all methods—including ours—perform inference using the original InternVL2.5-8B model.
Implementation details.
For EMLoC, we use LoRA with rank $8$ and train with $500$ iterations. The learning rate is set to $4\times 10^{-5}$ with cosine annealing. The number of calibration data is set to $64$ and $\lambda$ for LoRA correction is $3$. The other baselines use the same hyperparameters setting except we adjust the LoRA rank so that the number of trainable parameters of all baselines are the same.
Table 2: Applying EMLoC to 26B and 38B models. Note that a 5B-sized emulator is considered, and all experiments are conducted under a 24GB memory budget.
Size Method PMC-VQA WebSRC WC-VQA
26B Zero-shot 49.9 77.1 51.0
EMLoC 52.5 80.9 52.6
38B Zero-shot 52.5 79.0 53.6
EMLoC 57.0 82.1 56.8
Table 3: Applying EMLoC to NLP tasks. Following LORAM [45], a LLaMA2 13B is used as the full model. All experiments are conducted under the same budget.
MATHQA GSM8K
w/o FT 32.6 24.3
LORAM-RAND 33.8 27.2
LORAM-STRU 33.8 24.6
EMLoC 33.9 29.8
Table 4: Comparisons of emulator construction settings. Note that, LORAM [45] requires external data and additional continual pretraining for memory efficient finetuning. We validate that both row pruning and SVD can be applied for our emulator construction with supperior performances on various tasks using only downstream data.
Compression
Method
Add. Data for
compression
Require Add.
Cont. Pre-train
Overhead ($\downarrow$)
(GPU-hours)
PMC-VQA WebSRC WC-VQA
Row-pruning Yes Yes 214 51.0 78.7 43.6
No No 1.7 51.1 76.6 44.1
SVD Yes Yes 230 51.3 74.6 42.9
No No 0.3 51.6 79.6 46.2
4.2 Main Results
Main Results.
We evaluate EMLoC using InternVL2.5-8B as the base model under three realistic memory-constrained fine-tuning scenarios. These settings simulate user environments with training memory budgets equivalent to: fine-tuning a 4B model (denoted as $50\%$ compression ratio in the table), fine-tuning a 2B model ($25\%$), and fine-tuning a 2B model using QLoRA [9] quantization ($25\%$+QLoRA). As shown in Table 1, EMLoC consistently outperforms all baselines across both standard and domain-specific VQA benchmarks. The improvements are especially significant on specialized datasets, where adaptation to the target domain is essential. Notably, EMLoC achieves performance closest to the upper bound set by full-model fine-tuning.
These results also underscore the adaptability of our approach under varying memory constraints, maintaining strong performance even under the tightest budgets. In contrast to the common strategy of fine-tuning smaller models, EMLoC provides a more effective alternative by better leveraging the capacity of the full model. In addition, comparisons with UPop and Offsite-tuning demonstrate the advantages of our downstream-aware emulator construction. Unlike row-pruning or layer-dropping approaches that restrict architectural flexibility, our method tailors the emulator to downstream data without altering model structure, enabling more effective and generalizable adaptation.
4.3 Analysis
Scalability on Model Size.
We further evaluate EMLoC on larger-scale models, including the 26B and 38B variants of InternVL2.5. For this experiment, we construct a 5B-sized emulator, making the fine-tuning cost roughly equivalent to training a 5B model. We use LoRA rank 6 for the 26B model and rank 4 for the 38B model. As shown in Table 3, EMLoC consistently improves upon the zero-shot performance of these large models, demonstrating its effectiveness even at greater model scales. Importantly, all fine-tuning were conducted with less than 24GB of GPU memory—substantially lower than the 95GB typically required for directly fine-tuning the 38B model. This means that users capable of running inference can also perform fine-tuning using our method, without additional hardware requirements. These results underscore the practicality and scalability of EMLoC for resource-constrained users.
Comparison of Emulator Construction Settings.
We compare two emulator construction settings. The first is our proposed downstream-aware construction, in which the emulator is derived directly from task-specific data without any additional training. The second follows the approach adopted in LORAM, where the emulator is first constructed using general-purpose data, followed by continual pretraining. For general data, we use Laion-COCO [17]—a commonly used dataset for vision-language pretraining—and perform continual pretraining using 8 V100 GPUs.
As shown in Table 4, when row pruning is used as the emulator construction method, both settings achieve comparable downstream performance. However, the continual pretraining stage required in LORAM incurs substantial computational overhead, making it impractical for individual users. In contrast, when SVD is used for emulator construction, our downstream-aware method not only eliminates the need for continual pretraining but also achieves superior performance. Moreover, it incurs the lowest construction overhead, demonstrating the efficiency and effectiveness of our downstream-aware SVD.
Table 5: Ablation study on activation-aware SVD and LoRA correction.
Activation-Aware SVD LoRA Correction PMC-VQA WebSRC WC-VQA
51.0 74.4 44.7
51.5 79.0 45.8
51.2 74.4 44.8
51.6 79.6 46.2
Robustness Across Modalities.
To evaluate the generality of our approach beyond vision-language tasks, we apply EMLoC to two challenging NLP datasets, MATHQA and GSM8K. We follow the experimental setup of LORAM and use LLaMA 2 13B [36] as the base model. As shown in Table 3, our method achieves comparable performance on MATHQA and outperforms LORAM on GSM8K, demonstrating its effectiveness in language-only settings. These results confirm that EMLoC is not limited to multimodal models and can be readily extended to other modalities. Furthermore, our method retains its core advantages—minimal memory overhead and no need for continual pretraining—making it more accessible to users with limited computational resources.
Effectiveness of Transferring LoRA From Emulator to Original Model.
Table 6: Performance comparisons of EMLoC before and after transferring the LoRA modules from the emulator to the original model. Note that running inference using emulator does not fully utilize the capacity of the original model and thus is not desirable.
Ratio Method ChartQA DocVQA InfoVQA TextVQA PMC-VQA WebSRC WC-VQA
50% Emulator 57.1 62.3 29.3 15.0 40.9 42.1 30.7
EMLoC 84.6 92.3 69.9 78.8 52.3 85.2 48.8
25% Emulator 19.3 6.3 10.8 6.9 31.2 4.0 19.1
EMLoC 84.2 92.3 70.0 79.0 51.6 79.6 46.2
Emulator 14.8 4.5 10.5 7.3 32.2 4.3 19.0
25%+ QLoRA EMLoC 84.2 91.9 67.5 79.0 50.9 78.1 44.8
To better understand the role of the emulator and the effectiveness of our LoRA correction strategy, we compare model performance before and after transferring the fine-tuned LoRA modules from the emulator to the original model. As shown in Table 6, performance on the emulator itself is significantly lower the final model after LoRA transfer. This gap underscores the importance of using the full model for inference and justifies our design choice of transferring LoRA back.
Interestingly, despite the poor performance of the emulator, the LoRA modules trained on it improve the original model’s downstream performance once transferred. This indicates that the fine-tuning process successfully extracts domain-specific knowledge, which is embedded in the LoRA weights and effectively transferred to the original model. These results highlight the effectiveness of our emulator-based training.
Figure 4: Sensitivity analysis of $\lambda$ in the LoRA correction algorithm. We plot performance on WC-VQA under different $\lambda$ values.
Ablation Study.
We conduct ablation experiments to evaluate the contribution of two key components in the EMLoC framework: activation-aware SVD for downstream-aware emulator construction, and the LoRA correction algorithm for mitigating training-inference misalignment. As shown in Table 5, removing either component results in a drop in performance. This highlights the importance of preserving downstream-relevant knowledge during emulator construction, and the necessity of correcting for discrepancies introduced by training and inference on different models.
Impact of Hyperparameter $\lambda$.
We investigate the effect of the hyperparameter $\lambda$, which constrains the norm of the correction term $\Delta$ in our LoRA correction algorithm. As shown in Fig. 4, performance varies with different choices of $\lambda$. In particular, when no constraint is applied (i.e., $\lambda\rightarrow\infty$), performance degrades noticeably. We observe that, in such cases, the correction term may have a significantly larger norm than the original LoRA output. We hypothesize that correction term with large norm will dominate and distort the LoRA modules, ultimately harming downstream performance. In contrast, applying an appropriate constraint on the correction magnitude helps preserve the learned knowledge in the LoRA modules while still compensating for model misalignment.
5 Conclusion
We present EMLoC, a novel framework that makes fine-tuning large foundation models as memory-efficient as inference. By introducing a downstream-aware emulator constructed via activation-aware SVD and performing all training on this low-rank approximation, EMLoC enables individual users to fine-tune large models without exceeding the memory budget required for inference. Our LoRA correction algorithm further improve the performance by addressing the mismatch introduced by training and inference on different model bases. Extensive experiments across diverse datasets and modalities demonstrate that EMLoC not only outperforms other baselines but also scales to models as large as 38B parameters on commodity hardware. These results open up new opportunities for accessible and scalable model customization, empowering broader adoption of large language models in specialized, resource-constrained settings.